# Mobile Edge Intelligence for Large Language Models: A Contemporary Survey

### v1

@@ V-B Parameter-sharing LLM Caching and Delivery
Considering the limited storage capacities of edge devices and frequent model fine-tuning, LLMs of interest should be delivered within edge networks from their location to where they are needed promptly. Moreover, considering RAG, external knowledge sources should also be cached at the network edge, ensuring timely acquisition of the data/knowledge once LLM applications need. The model/data delivery can be carried out in wired backhaul or radio access networks. The caching and delivery of LLMs must take advantage of the fact that parameter blocks can be shared among various downstream LLMs [160, 112] and even reused within the same LLM [38]. This unique characteristic enables efficient edge LLM caching and low-latency LLM delivery by reducing caching and delivery costs of duplicate LLM parameter blocks, as we will present in VI. To enable fast model delivery, MEI4LLM can construct a lookup table, assigning names for LLM parameter blocks to facilitate content search and management, following the principle of information-centric networking. By doing so, the MEI4LLM paradigm places LLMs at appropriate sites, retrieves LLMs of interest from nearby edge servers, and enables routing/multicasting of LLM parameter blocks to mobile users.
@@ V-C Distributed LLM Training (Fine-tuning)
It is foreseen that 6G MEI systems can efficiently fine-tune LLMs to adapt to local environments. Edge LLM fine-tuning can be triggered whenever the inference accuracy decreases or after a certain period when local environments have changed. For example, LLM-empowered virtual assistants should be fine-tuned regularly to better adapt to new trends in news media, top local restaurants, and popular attractions, resulting in improved decision-making and interactions with users. LLM-empowered mobile health applications should be personalized to provide better predictions and health or fitness suggestions.
In next-generation mobile networks, edge training for LLMs must answer two questions: 1) how to preserve user privacy and data ownership, and 2) how to support large-scale model training through the collaboration of edge nodes. To enhance the data privacy of users, federated learning (FL) and split learning (SL) serve as two promising distributed learning frameworks to implement at the network edge. Specifically, FL allows edge devices to train models locally while only sharing model parameters with an edge server for aggregation, thus harnessing collective intelligence without sharing personal data. Alternatively, SL and its variant split federated learning (SFL), can be implemented to enable device-server co-training without sharing local raw data, which is particularly suitable for large-scale LLM fine-tuning for edge devices [161], as model splitting allows for workload balancing across diverse edge nodes. To effectively support the intensive training, various resource-efficient training techniques, as detailed in Section III, can be combined with FL or SL. These discussions will be provided in Section VII.
@@ V-D Distributed LLM Inference
To accommodate for resource-intensive LLMs, edge servers and edge devices must perform distributed inference for LLMs in a concerted manner, depending on the communication-computing workload and privacy requirements. There are different ways for edge inference. On-server inference requires users to upload raw data to the server. This approach eliminates the computing burden on edge devices while potentially violating users’ privacy requirements. For instance, multimodal LLMs may collect sensitive audio and video data in home environments, where users are often reluctant to share. Conversely, on-device inference preserves privacy and eliminates communication costs while imposing an intensive computing workload on edge devices. Split inference, a key AI inference framework in 3GPP 5G technical specifications [79], sits in between, with edge devices and servers holding parts of the AI models. Split inference involves uploading features from edge devices to edge servers for co-inference. To facilitate LLM inference, MEI4LLM can tailor consumer services by appropriately selecting from these schemes based on communication-computing resource status and privacy requirements, as elaborated in Section VIII.
@@ VI Edge Caching and Delivery for LLMs
Edge LLM caching and delivery play indispensable roles in both the training and inference of LLMs, serving as cornerstones of edge LLM deployment. For this reason, we begin with discussions on edge caching and delivery and then present edge training and inference in the subsequent sections. Compared with conventional edge service/content caching and delivery, the main distinction of edge LLM caching and delivery is the exploitation of parameter shareability, which is commonly observed in LLMs, aiming to increase storage and communication efficiency over edge networks. While parameter shareability can exist in traditional DNNs, it is much more prevalent and significant in LLMs due to the widespread adoption of PEFT techniques, requiring our special design attention. In what follows, we present the techniques arising from exploiting this characteristic of LLMs. The related works for efficient edge LLM caching and delivery are outlined in Table III for readers’ convenience.
TABLE III: Summary of related works for edge LLM caching and delivery.
@@ VI-A Edge LLM Caching
Edge model caching can realize low model downloading latency by distributing AI models to wireless edge servers in advance. Distinct from service placement for computation offloading, edge model caching focuses on caching AI models for downloading from edge servers to end users. The design of AI model caching aims to provide models to more users within their quality-of-service (QoS) requirements [160, 165]. This paradigm enables users to fetch AI models from edge servers directly instead of accessing remote cloud data centers, which incurs excessive downloading latency [164, 162]. However, implementing edge LLM caching presents several challenges: 1) Limited storage capacity for LLM caching: Service providers aim to place as many popular LLMs as possible on edge servers to enhance the cache hit ratio and decrease model downloading latency for users. Nevertheless, the immense size of LLMs presents a significant challenge for their storage on edge servers; 2) High LLM edge cache (re)placement costs: Over time, previously cached LLMs may no longer align with the changing user requests. To address this, service providers may replace LLMs on edge servers to better accommodate up-to-date requests. However, the placement of these large-scale models leads to considerable communication overhead and imposes a substantial burden on mobile backhaul networks. In what follows, parameter-sharing model caching is presented to address the above challenges.
@@ VI-A1 Parameter-sharing LLM caching
Parameter-sharing model caching can be adopted to improve storage and transmission efficiency at the network edge. As discussed in Section III, PEFT, such as LoRA, is widely adopted for adapting LLMs to downstream tasks. In LoRA, the pre-trained LLM parameters are frozen, and only the newly introduced parameters are trainable, typically accounting for less than $1\%$ of the original LLM parameters. Therefore, most parameters of various LLMs fine-tuned with LoRA for downstream tasks are shared from pre-trained LLMs, which should be leveraged to enhance caching efficiency significantly. Take LoRA and GPT-2 as an example. Fig. 7 demonstrates that the inference performance almost remains unchanged even when 99.97% of parameters in the GPT-2 large model are frozen parameters from the pre-trained GPT-2 large. Based on this observation, in [162], we propose an AI model placement strategy, called TrimCaching, to maximize the cache hit ratio under server storage capacity and service latency constraints by exploiting the parameter-sharing properties of AI models, particularly LLMs. In the TrimCaching framework, only one copy of the shared parameter blocks across LLMs is cached on one edge server, thereby improving storage efficiency, as illustrated in Fig. 8. Compared with the Independent Caching strategy for edge LLM caching [166, 167], which does not consider the parameter sharing across LLMs, the TrimCaching strategy can significantly improve cache hit ratio as shown in Fig. 9.
@@
Figure 9: Cache hit ratio v.s edge server capacity. We employ LoRA to fine-tune GPT-2 models for 100 downstream tasks across 40 users in a scenario with 12 edge servers. A cache hit occurs if the requested GPT-2 model can be served by any edge server in the edge networks within the E2E latency requirements of end users. The other parameter settings can be referred to [162]. It is noted that the capacity of edge servers in the real world is much larger than 5 GB. However, the number of downstream tasks for LLMs is also more than 100 in real life. Due to the limited computing resources, the number of downstream tasks and the edge server capacity are reduced proportionally, which does not affect the observations.
@@ VI-A2 Edge LLM cache replacement
Since the popularity of models can evolve over time, another fundamental research problem in edge caching is LLM replacement. By replacing outdated content with new data, edge servers can continuously refresh their caches with new content to satisfy the ever-changing user requests [169]. The two most classic replacement strategies are the recency-based and the frequency-based strategies, which remove the least recently used (LRU) objects and least frequently used (LFU) objects and then replace them with updated content. However, such strategies fail to consider the shared parameter blocks across LLMs and the cooperative caching among different edge nodes. One direction to improve the replacement performance is to re-conduct the centralized proactive caching, say the scheme in [160], after a certain period. However, under high user mobility, such methods may involve high system complexity and high communication costs for transferring LLMs. Another direction is to develop distributed algorithms, often based on the Markov Decision Process or reinforcement learning, to make replacement decisions without knowing the full information of other edge nodes [170].
@@ VI-A3 Edge caching for RAG
As mentioned in Section II-A4, integrating LLMs and RAG, which enables LLMs to retrieve relevant data from external knowledge sources, is essential for generating reliable and up-to-date responses without retraining/fine-tuning. However, since retrieving the information from the remote cloud can be time-consuming, the most popular external knowledge should be cached at the network edge to enable LLMs to fetch the most factual, accurate, and up-to-date content. Intriguingly, this edge caching problem naturally differs from traditional ones as the caching should be optimized by considering the training status or internal knowledge of LLMs. Specifically, if LLMs can already memorize or easily infer certain content, such knowledge can be removed from the external knowledge sources to save storage space at the network edge, thereby enhancing storage efficiency for caching external knowledge sources. Besides, edge servers can cache specific external knowledge sources frequently requested by their associated users, which can significantly increase the QoS for responses to users in different regions. On the other hand, edge caching for RAG is tightly coupled with on-the-fly LLM fine-tuning, i.e., we need to choose which data to fine-tune outdated LLMs and which data to cache for RAG on edge servers. With this in mind, it is worth investigating a joint LLM fine-tuning and knowledge source caching problem under the context of RAG to enhance the reliability of LLMs under latency constraints, taking into account the data retrieval latency from external knowledge sources at the edge/cloud servers and the fine-tuning costs by feeding new data to LLMs.
@@ VI-B2 Parameter-sharing wireless model downloading
Parameter-sharing LLM delivery/downloading can also be integrated with various LLM compression techniques. Take quantization as an example; by compressing LLM weights in low bitwidth, LLM downloading can be accomplished with low latency [171, 164]. However, as illustrated in Section III-A, setting a uniform quantization bitwidth in LLMs degrades the inference performance significantly since not all weights contribute equally to the final outputs [90]. Therefore, weights in LLMs can be assigned with non-uniform bitwidths according to weight importance [172]. For example, weights corresponding to larger activations or higher quantization errors can be quantized in high bitwidth before downloading, which can preserve the performance of the quantized LLMs [90, 173, 172]. Moreover, quantization should also be optimized by considering the popularity/shareability of parameter blocks across various downstream LLMs. By integrating LLM compression methods, the design for LLM delivery/downloading by jointly considering wireless channel conditions, weight importance, parameter block shareability, and model performance is worth further investigation.
@@ VII Edge Training for LLMs
Edge training performs model training at the network edge to extract intelligence from data sources. The main distinction of edge LLM training from traditional edge training lies in the large scale of AI models, which can be too large to fit into an edge server, as well as the optimization of PEFT over wireless networks. As shown in Fig. 12, we elaborate on edge LLM training through four categories: centralized edge learning, federated edge learning, split learning, and hierarchical collaborative learning. For readers’ convenience, the related works on edge LLM training are summarized in Table IV.
(a) Centralized edge learning.
(b) Federated edge learning.
(c) Split learning.
(d) Hierarchical collaborative learning.
Figure 12: Edge training frameworks for LLMs. This figure presents the associated challenges and corresponding techniques to address them in each sub-figure. In Fig. 12(a), edge devices upload raw data to the edge server for global LLM training. In Fig. 12(b), FL is adopted, where the edge server aggregates the locally trained LLMs from edge devices. In Fig. 12(c), an LLM is partitioned into two sub-models, with the split learning process executed by exchanging intermediate data between the edge server and the edge device. At last, hierarchical collaborative learning can be categorized into cloud-end, cloud-edge, and cloud-edge-end collaborative learning. We take cloud-edge collaborative learning as an example in Fig. 12(d), where edge devices upload the data to edge servers for training, and the cloud aggregates different edge-trained LLMs.
TABLE IV: Summary of related works on edge LLM training.
@@ VII-A1 Scaling-up LLM training
Scaling-up training is essential for LLM training in centralized edge training. Due to memory and computational constraints, training large-scale LLMs on a single GPU is highly challenging. Therefore, in the context of centralized edge training, scaling up with distributed computing and memory resources, such as multiple GPUs on edge servers, can accelerate the training process and make it feasible to train LLMs in edge networks. Scaling-up LLM training can be categorized into two main approaches: parallel training and GPU memory optimization, which are elaborated below.
Parallel training for LLMs: Considering the extreme training workload of LLMs, parallel computing must be employed to leverage resources across edge nodes to split LLM training to reduce training latency and share required memory space. The three most prominent parallel computing strategies are data parallelism, pipeline parallelism, and tensor parallelism [196, 197]: 1) Data parallelism enables replicating models across different processors (processors can refer to multiple GPUs in one server or different edge servers), with the uploaded dataset being shuffled and distributed among the processors. After each processor finishes the forward and backward propagation with the corresponding model replica in parallel, all the gradients are aggregated for synchronization to update the model parameters [175]. 2) In pipeline parallelism, models are divided into different sub-models; each processor owns one sub-model. A mini-batch of training samples is further divided into micro-batches in each training iteration. The pipeline training process can be executed by feeding micro-batches into the sub-model, and different processors can update the sub-models with different micro-batches at the same time [176]. 3) In tensor parallelism, the parameters and computations of each layer are divided across multiple processors for separate processing by multiple processors in parallel. For example, considering Transformers, the authors in [177] split per attention head parameters in self-attention modules across multiple processors such that the matrix multiplication corresponding to each attention head is stored locally on one processor. To leverage the advantages of these three parallel computing paradigms, in [178], Narayanan et al. combine tensor, pipeline, and data parallelism to efficiently train LLMs by harnessing a massive number of distributed GPUs.
@@
VII-A2 Resource management for scaling-up LLM training
Different from cloud-based approaches, resource management for parallel computing of LLMs plays an essential role at the network edge. Although the aforementioned cloud-based methods can be applied to interconnected edge servers, the heterogeneous communication and computing capabilities of edge servers must be carefully considered for system optimization, which is normally oversimplified in large-scale GPU clusters. For example, in pipeline parallelism, when placing sub-models to multiple edge servers for LLM training, the intermediate smashed data should be exchanged among edge servers over significantly heterogeneous wired/wireless communication links. In this regard, model splitting and placement should be optimized judiciously based on the computing-communication resource constraints. In fact, similar problems have been considered for multi-hop split inference [199, 200], where the model splitting and/or placement is often mapped to graph construction problems, which can be solved by shortest path problem [200] or min-cut problem [199] considering the communication-computing delay with different splitting and placement choices. Unfortunately, it is noted that these problems are not directly applicable to parallel training of large-scale models, such as pipeline parallelism and tensor parallelism under edge networks. For instance, unlike multi-hop inference problems that often handle one input data at a time, pipeline parallelism for model training considers processing a batch of training data samples, which can be divided into multiple micro-batches as inputs to form the pipeline process. For this reason, the end-to-end delay consists of the delay of processing one micro-batch of data samples and the pipeline delay of processing all micro-batches in the entire batch. These characteristics set apart parallel training for large-scale models from inference problems in edge computing networks. Besides, by exploiting the special structure of LLMs, mostly relying on Transformers, the workload splitting problem might be simplified because it consists of repeated Transformer blocks with the same size of intermediate data output. Furthermore, model splitting, tensor splitting, and sub-model/layers placement can be jointly optimized with radio resource management problems, such as spectrum allocation and data routing [201], to mitigate network congestion with reduced transmission delay.

### v2

@@ V-B Parameter-sharing LLM Caching and Delivery
Given the limited storage capacities of edge devices and real-time model fine-tuning in changing environments, LLMs must be delivered over mobile networks frequently to support subsequent usage. Since parameter blocks can be shared among various downstream LLMs [186, 153] and even reused within the same LLM [70], LLM caching and delivery schemes must take advantage of this fact to reduce caching and delivery costs of shared LLM parameter blocks. Besides, to enable real-time model delivery, MEI4LLM can construct a lookup table, assigning names for LLM parameter blocks to facilitate content search and management, following the principle of information-centric networking [187, 188]. By doing so, the MEI4LLM paradigm places LLMs at appropriate sites, retrieves LLMs of interest from nearby edge servers, and enables routing/multicasting of LLM parameter blocks to mobile users. Details will be further discussed in Section VI.
@@ V-C Distributed LLM Training (Fine-tuning)
It is foreseen that 6G MEI systems can efficiently fine-tune LLMs to adapt to local environments. Edge LLM fine-tuning can be triggered whenever the inference accuracy decreases or after a certain period when local environments have changed. For example, LLM-empowered virtual assistants should be fine-tuned regularly to better adapt to new trends in news media, top local restaurants, and popular attractions, resulting in improved decision-making and interactions with users [189, 190]. LLM-empowered mobile health applications should be personalized to provide better predictions and health or fitness suggestions [191].
In next-generation mobile networks, edge training for LLMs must answer two questions: 1) how to preserve user privacy and data ownership, and 2) how to support large-scale model training through the collaboration of edge nodes. To enhance the data privacy of users, federated learning (FL) and split learning (SL) serve as two promising distributed learning frameworks to implement at the network edge [192]. Specifically, FL allows edge devices to train models locally while only sharing model parameters with an edge server for aggregation, thus harnessing collective intelligence without sharing personal data [193]. Alternatively, SL and its variant split federated learning (SFL), can be implemented to enable device-server co-training without sharing local raw data, which is particularly suitable for large-scale LLM fine-tuning for edge devices [194], as model splitting allows for workload balancing across diverse edge nodes. To effectively support the intensive training, various resource-efficient training techniques, as detailed in Section IV, can be combined with FL or SL. These discussions will be provided in Section VII.
In addition, considering RAG [195, 196], external knowledge sources should also be cached at the network edge, ensuring timely acquisition of the data/knowledge for an LLM. However, RAG involves extra storage costs and delivery latency for retrieving external knowledge. In comparison, fine-tuning an LLM consumes extra computing power while eliminating the need for external databases. In this regard, an important research problem is determining data selection for either LLM training (fine-tuning) or RAG in MEI systems.
@@ V-D Distributed LLM Inference
To accommodate for resource-intensive LLMs, edge servers and edge devices must perform distributed inference for LLMs in a concerted manner, depending on the communication-computing workload and privacy requirements. There are different ways for edge inference. On-server inference requires users to upload raw data to the server. This approach eliminates the computing burden on edge devices while potentially violating users’ privacy requirements. For instance, multimodal LLMs may collect sensitive audio and video data in home environments, where users are often reluctant to share. Conversely, on-device inference preserves privacy and eliminates communication costs while imposing an intensive computing workload on edge devices. Split inference sits in between, with edge devices and servers holding parts of the AI models, which has been mentioned in 3GPP 5G technical specifications. Split inference involves uploading features from edge devices to edge servers for co-inference. To facilitate LLM inference, MEI4LLM can tailor consumer services by appropriately selecting from these schemes based on communication-computing resource status and privacy requirements, as elaborated in Section VIII.
V-E Lessons Learned for MEI4LLM
Clearly, MEI4LLM is nothing but a special case of MEI. However, the need to train and deploy LLMs at the edge serves as a key stimulus for the evolution of MEI towards its next stage. Specifically, the extreme resource demands of LLMs necessitate a fundamental rethinking of traditional MEI. For this reason, MEI4LLM must feature the following characteristics: 1) the native support of model splitting and parallel training/inference across interconnected edge nodes to facilitate the deployment of large-scale models 2) the integrated design of wireless communications and resource-efficient LLM training/inference techniques such as parameter-efficient fine-tuning and token (representation) reduction to make LLM deployment cost-effective. The remainder of this survey paper will discuss these two critical aspects in detail.
@@ VI Edge Caching and Delivery for LLMs
Edge LLM caching and delivery play indispensable roles in both the training and inference of LLMs, serving as cornerstones of edge LLM deployment. For this reason, we begin with discussions on edge caching and delivery and then present edge training and inference in the subsequent sections. Compared with conventional edge service/content caching and delivery, the main distinction of edge LLM caching and delivery is the exploitation of parameter shareability, which is commonly observed in LLMs, aiming to increase storage and communication efficiency over edge networks. While parameter shareability can exist in traditional DNNs, it is much more prevalent and significant in LLMs due to the widespread adoption of PEFT techniques, requiring our special design attention. In what follows, this section presents the techniques that leverage this characteristic of LLMs for optimizing edge caching and delivery. The related works for efficient edge LLM caching and delivery are outlined in Table IV for readers’ convenience.
TABLE IV: Summary of related works for edge LLM caching and delivery.
@@ VI-A Edge LLM Caching
Edge model caching can realize low model downloading latency by distributing AI models to wireless edge servers in advance. Distinct from service placement for computation offloading, edge model caching focuses on caching AI models for downloading from edge servers to end users [186, 200]. This paradigm enables users to fetch AI models from edge servers directly instead of accessing remote cloud data centers, which incurs excessive downloading latency [199, 197]. However, implementing edge LLM caching presents several challenges: 1) Limited storage capacity for LLM caching: Service providers aim to place as many popular LLMs as possible on edge servers to enhance the cache hit ratio and decrease model downloading latency for users. Nevertheless, the immense size of LLMs presents a significant challenge for their storage on edge servers; 2) High LLM edge cache (re)placement costs: Over time, previously cached LLMs may no longer align with the changing user requests. To address this, service providers may replace LLMs on edge servers to better accommodate up-to-date requests, imposing a substantial burden on mobile backhaul networks. In what follows, parameter-sharing model caching is presented to address the above challenges.
@@ VI-A1 Parameter-sharing LLM caching
Parameter-sharing model caching can be adopted to improve storage and transmission efficiency at the network edge. As discussed in Section IV, PEFT, such as LoRA, is widely adopted for adapting LLMs to downstream tasks. In LoRA, the pre-trained LLM parameters are frozen, and only the newly introduced parameters are trainable, typically accounting for less than 1% of the original LLM parameters. Therefore, most parameters of various LLMs fine-tuned with LoRA for downstream tasks are shared from pre-trained LLMs, which should be leveraged to enhance caching efficiency significantly. Take LoRA and GPT-2 as an example. Fig. 10 demonstrates that the inference performance almost remains unchanged even when 99.97% of parameters in the GPT-2 large model are frozen parameters from the pre-trained GPT-2 large. Based on this observation, in [197], we propose an AI model placement strategy, called TrimCaching, to maximize the cache hit ratio under server storage capacity and service latency constraints by exploiting the parameter-sharing properties of AI models, particularly LLMs. In the TrimCaching framework, only one copy of the shared parameter blocks across LLMs is cached on one edge server, thereby improving storage efficiency, as illustrated in Fig. 11. Compared with the Independent Caching strategies for edge LLM caching [201, 202], which do not consider parameter sharing across LLMs, the TrimCaching strategy can significantly improve the cache hit ratio as shown in Fig. 12. In other words, even with storage-limited edge servers, a large number of popular LLMs can still be cached on network edge servers under the TrimCaching framework, thereby considerably reducing LLM downloading latency compared with fetching them from the cloud.
@@
Figure 12: Cache hit ratio v.s edge server capacity. A cache hit occurs if the requested GPT-2 model can be served by any edge server in the edge networks within the E2E latency requirements of end users. The GPT-2 models, including GPT-2 small, medium, and large, are fine-tuned with LoRA for parameter sharing in 100 downstream tasks. 40 users and 12 edge servers are located in a square area of 1 $\text{km}^{\text{2}}$. The coverage radius of edge servers is 275 m. The other parameter settings can be referred to [197].
@@ VI-A2 Edge LLM cache replacement
Since the popularity of models can evolve over time, another fundamental research problem in edge caching is LLM replacement. By replacing outdated content with new data, edge servers can continuously refresh their caches with new content to satisfy ever-changing user requests [204]. The two most classic replacement strategies are the recency-based and the frequency-based strategies, which remove the least recently used (LRU) objects and least frequently used (LFU) objects and then replace them with updated content. However, such naive strategies fail to exploit the cooperative caching among edge nodes and parameter shareability among LLMs. Two directions can be further explored to improve the performance. First, shared parameter blocks among LLMs and cooperation among neighboring edge servers can be considered to achieve parameter replacement with minimal backhaul transmission costs. Second, model replacement should be optimized with inference accuracy in mind, as LLMs are often more effective in general tasks than specific ones. This suggests that users may request an LLM that closely matches their needs, i.e., achieving required prediction accuracy rather than a specific LLM, thereby leading to less frequent replacement.
To implement the replacement policies, one direction is centralized proactive caching, say the scheme in [186], after a certain period. However, under high user mobility, such methods may involve high system complexity and communication costs for transferring LLMs. The other direction is to develop distributed algorithms, often based on the Markov Decision Process or reinforcement learning, to make replacement decisions without knowing the full information of other edge nodes [205]. Both of the directions are worth exploring further.
@@ VI-A3 Edge caching for RAG
As mentioned in Section III-A4, integrating LLMs and RAG, which enables LLMs to retrieve relevant data from external knowledge sources, is essential for generating reliable and up-to-date responses without retraining/fine-tuning. However, since retrieving the information from the remote cloud can be time-consuming, the most popular external knowledge should be cached at the network edge to enable LLMs to fetch the most factual, accurate, and up-to-date content. Intriguingly, this edge caching problem naturally differs from traditional ones as the caching should be optimized by considering the training status or internal knowledge of LLMs. Specifically, if LLMs can already memorize or easily infer certain content, such knowledge can be removed from the external knowledge sources to save storage space at the network edge, thereby enhancing storage efficiency for caching external knowledge sources. Besides, edge servers can cache specific external knowledge sources frequently requested by their associated users, which can significantly increase the QoS for responses to users in different regions. On the other hand, edge caching for RAG is tightly coupled with on-the-fly LLM fine-tuning, i.e., we need to choose which data to fine-tune outdated LLMs and which data to cache for RAG on edge servers. With this in mind, it is worth investigating a joint LLM fine-tuning and knowledge source caching problem under the context of RAG to enhance the reliability of LLMs under latency constraints, taking into account the data retrieval latency from external knowledge sources at the edge/cloud servers and the fine-tuning costs by feeding new data to LLMs.
@@ VI-B2 Parameter-sharing wireless model downloading
Parameter-sharing LLM delivery/downloading can also be integrated with various LLM compression techniques. Take quantization as an example; by compressing LLM weights in low bitwidth, LLM downloading can be accomplished with low latency [206, 199]. However, as illustrated in Section IV-A, setting a uniform quantization bitwidth in LLMs degrades the inference performance significantly since not all weights contribute equally to the final outputs [131]. Therefore, weights in LLMs can be assigned with non-uniform bitwidths according to weight importance [207]. For example, weights corresponding to larger activations or higher quantization errors can be quantized in high bitwidth before downloading, which can preserve the performance of the quantized LLMs [131, 208, 207]. Moreover, quantization should also be optimized by considering the popularity/shareability of parameter blocks across various downstream LLMs. By integrating LLM compression methods, the design for LLM delivery/downloading by jointly considering wireless channel conditions, weight importance, parameter block shareability, and model performance is worth further investigation.
@@ VII Edge Training for LLMs
Edge training performs model training at the network edge to extract intelligence from data sources. The main distinction of edge LLM training from traditional edge training lies in the large scale of AI models, which can be too large to fit into an edge server, as well as the optimization of PEFT over wireless networks. As shown in Fig. LABEL:fig:edge_learning, this section elaborates on edge LLM training through four categories: centralized edge learning, federated edge learning, split learning, and hierarchical collaborative learning. For readers’ convenience, the related works on edge LLM training are summarized in Table V.
TABLE V: Summary of related works on edge LLM training.
@@ VII-A1 Scaling-up LLM training
Scaling-up training is essential for LLM training in centralized edge training. Due to memory and computing constraints, training large-scale LLMs on a single GPU is highly challenging. Therefore, leveraging distributed computing and memory resources, such as multiple GPUs on edge servers, come to the rescue. Scaling-up LLM training can be categorized into two main approaches: parallel training and GPU memory optimization, which are elaborated below.
Parallel training for LLMs: Considering the extreme training workload of LLMs, parallel computing must be employed to leverage resources across edge nodes to split LLM training to reduce training latency and share required memory space. The three most prominent parallel computing strategies are data parallelism, pipeline parallelism, and tensor parallelism [231, 232]: 1) Data parallelism enables replicating models across different processors (processors can refer to multiple GPUs in one server or different edge servers), with the uploaded dataset being shuffled and distributed among the processors. After each processor finishes the forward and backward propagation with the corresponding model replica in parallel, all the gradients are aggregated for synchronization to update the model parameters [210]. 2) In pipeline parallelism, models are divided into different sub-models; each processor owns one sub-model. A mini-batch of training samples is further divided into micro-batches in each training iteration. The pipeline training process can be executed by feeding micro-batches into the sub-model, and different processors can update the sub-models with different micro-batches at the same time [211].
Figure 18: Per-round latency for split edge learning based on the number of edge servers participating in multi-hop split edge learning. We train a LLaMA-2 model on the E2E dataset [233] to evaluate performance, assuming a wireless communication data rate of 200 Mbps between two edge servers. Per-round latency refers to the time taken to train one mini-batch of 128 samples.
3) In tensor parallelism, the parameters and computations of each layer are divided across multiple processors for separate processing by multiple processors in parallel. For example, considering Transformers, the authors in [212] split per attention head parameters in self-attention modules across multiple processors such that the matrix multiplication corresponding to each attention head is stored locally on one processor. To leverage the advantages of these three parallel computing paradigms, in [213], Narayanan et al. combine tensor, pipeline, and data parallelism to efficiently train LLMs by harnessing a massive number of distributed GPUs.
@@
VII-A2 Parallel LLM training at the network edge
Parallel computing for LLM training can be extended to wireless edge networks. As illustrated by our simulations in Fig. 18, pipeline parallelism significantly reduces end-to-end training latency by splitting the model into multiple sub-models and placing them on edge servers appropriately. Different from cloud-based approaches, parallel computing of LLMs encounters a more significant communication bottleneck at the network edge. Specifically, there exist heterogeneous communication and computing capabilities of edge servers. This situation, however, is normally oversimplified in large-scale GPU clusters. Since high-dimensional smashed data should be exchanged among edge servers via wired/wireless links, model splitting and placement should be optimized judiciously under computing-communication resource constraints.
Similar problems have been considered for multi-hop split inference [235, 236], where the model splitting and/or placement is often mapped to a graph, which can be further mapped to shortest path problem [236] or min-cut problem [235]. Unfortunately, these problems are not directly applicable to parallel training of large-scale models, such as pipeline parallelism and tensor parallelism under edge networks. For instance, unlike multi-hop inference problems that often handle one input data at a time, pipeline parallelism for model training considers processing a batch of training data samples, which can be divided into multiple micro-batches as inputs to form the pipeline process. These characteristics set apart parallel training for large-scale models from inference problems in edge computing networks. Besides, by exploiting the special structure of LLMs, mostly relying on Transformers, the workload splitting problem might be simplified because it consists of repeated Transformer blocks with the same size of intermediate data output. Furthermore, model splitting, tensor splitting, and sub-model/layers placement can be jointly optimized with radio resource management problems, such as spectrum allocation and data routing [237], to mitigate network congestion with reduced transmission delay.